OpenStudy (rockycatsrule):

What is the exact volume of a sphere with a diameter of 18 cm? cm3

OpenStudy (rockycatsrule):

d=2r=2·18=36

OpenStudy (rockycatsrule):

so 36 is my answer

OpenStudy (rockycatsrule):

@KamiBug

OpenStudy (anonymous):

(4/3)pi*r^3 = volume of sphere

OpenStudy (rockycatsrule):

V≈24429.02

OpenStudy (anonymous):

If you have a diameter of 18cm's, what's your radius?

OpenStudy (rockycatsrule):

r≈1.63

OpenStudy (anonymous):

the radius is (1/2) the diameter

OpenStudy (rockycatsrule):

duh

OpenStudy (anonymous):

So you have (4/3)*pi*9^3

OpenStudy (anonymous):

9^3 = 729 (4/3)*pi*729